the length of a rectangle is 2cm more than 3 times the width. the area is 16Cm.find the length and the width.let x=be the width
Answer by checkley79(3341)

L=3X+2
AREA=LX
16=(3X+2)X
16=X^2+2X
3X^2+2X-16=0
(X+8)X-2)=0
X-2=0
X=2 CM. ANS. FOR THE WIDTH
PROOF:
16=(3*2+2)2
16=(6+2)2
16=8*2
16=16